# Towards automated formal security analysis of SAML V2.0 Web Browser SSO standard — the POST/Artifact use case

### v1

@@ Abstract
Single Sign-On (SSO) protocols streamline user authentication with a unified login for multiple online services, improving usability and security. One of the most common SSO protocol frameworks — the Security Assertion Markup Language V2.0 (SAML) Web SSO Profile — has been in use for more than two decades, primarily in government, education and enterprise environments. Despite its mission-critical nature, only certain deployments and configurations of the Web SSO Profile have been formally analyzed. This paper attempts to bridge this gap by perforing a comprehensive formal security analysis of the SAML V2.0 SP-initiated SSO with POST/Artifact Bindings use case. Rather than focusing on a specific deployment and configuration, we closely follow the specification with the goal of capturing many different deployments allowed by the standard. Modeling and analysis is performed using Tamarin prover — state-of-the-art tool for automated verification of security protocols in the symbolic model of cryptography. Technically, we build a meta-model of the use case that we instantiate to eight different protocol variants. Using the Tamarin prover, we formally verify a number of critical security properties for those protocol variants, while identifying certain drawbacks and potential vulnerabilities.
@@
security protocols, formal proof, formal verification, SAML, security analysis, Single Sing-On, SSO, Tamarin prover
@@ I Introduction
Increasing number of online services are requiring user authentication, with passwords still being the prevalent authentication method. Accessibility issues related to unique strong passwords naturally lead to password reuse, which, in turn, makes services vulnerable to various attacks [1, 2], degrading the overall security on the Internet. Simultaneously, the burden on service providers to deploy, maintain and safeguard the authentication infrastructure further compounds the complexity. In response to these challenges, the concept of Web Single Sign-On (SSO) emerged at start of this millennium. SSO protocols enable exchange of authentication and authorization data between trusted identity providers (IdPs) and service providers (SPs), enabling users to authenticate once at an IdP and access services across multiple SPs.
@@
SSO frameworks ease the burden of authentication on SPs, but also impose stricter security requirements for IdPs and the SSO framework itself, as they are a single point of failure — a compromised IdP or an SSO system weakness can lead to compromises of sensitive user data across different SPs. For example, in the Sunburst [5, 6] supply chain attacks, after the initial compromise, the attackers obtained trusted signing certificate used for signing SAML and used it to access all services within the targeted enterprise.
@@
There have been numerous efforts in empirical security analysis of OAuth 2.0 [7, 8, 9, 10, 11, 12], OIDC [13] and SAML [14, 15, 16, 17], resulting in improved protocol implementations and specifications. Whilst the empirical security analysis can identity flaws and improve security, it can not guarantee the absence of weaknesses. On the other hand, formal security analysis can offer valid proofs, sometimes machine verified or even machine generated, that the mathematical model of the protocol holds desired security properties. Of course, the security guarantees provided by a deployed system depend on the faithfulness of the mathematical model to the specification and implementation.
@@
Despite its mission-critical nature and wide adoption, SAML V2.0 has received very little attention from the formal methods community, with the work by Armando et. al [18, 19, 20] being the only attempt to formally analyze systems based on the SAML framework. Their research, however, is exclusively focused on the so called Redirect/POST use case which is only one of several patterns for building Web SSO with SAML (albeit the most common one in use today). Furthermore, Armando et al. only considered one specific deployment configuration used by Google, while SAML standard allows for multiple protocol variants and features. There are no attempts at formal analysis of other use cases, let alone to model the entire specification.
@@
One of the obstacles for formal analysis is that SAML is not itself a security protocol — it is a set of building blocks that can be combined, configured and implemented in many different ways to build the protocol that achieves the desired SSO functionality. Even when we fix the specific SSO use case (e.g., the Redirect/POST use case is this paper), there are at least two additional sources of flexibility — optional features and under-specified mechanisms. While such flexibility is sometimes purely functional, there are cases where both optional features and under specified mechanisms do have an influence on the resulting security properties.
@@
In this paper, we focus on the SP-initiated SSO with POST/Artifact bindings use case and attempt to model the specification, accounting for the flexibility allowed by the standard. We develop a straightforward methodology that, first, identifies sensitive optional features and sensitive under-specifies mechanisms. Such features and mechanisms can be instantiated in several ways with each instantiation resulting in what we call a protocol variant. An example of optional features are digital signatures of authentication requests, which are recommended (but not required) by the SAML standard. Finally, we perform automated analysis of protocol variants yielding security proofs or falsifications of desired properties.
@@
There are multiple benefits to this approach. Most importantly, we can use automated formal analysis tools to examine the exact impact of each feature on the overall security guarantees provided by the protocol, as SAML standards only gives vague low level guidelines. Using the example of optional signatures, the standard [21] says ”The SAML request MAY be signed, which provides both authentication of the requester and message integrity.” — it is unclear what exactly (if anything) can go wrong in systems configured not to require signed authentication requests. Our analysis gives an answer to this question — for the POST/Artifact use case, adding signatures to authentication requests does not have a significant influence on the security of the use case in the symbolic model.
@@
The analysis presented in this paper, while significantly more comprehensive than the related work [18, 19, 20], still fails short of fully modelling the use case since we only consider three degrees of freedom, and model only two options for each. However, this is the first analysis of an important and widely used SAML Web SSO use case (three EU countries, support for this use case as part of their eID services [22]). Furthermore, the result of the formal analysis validate the methodology, demonstrating that protocol variants do indeed have different security features.
@@
In order to perform the analysis we needed to overcome a number of technical challenges related both to SAML itself and modeling with the Tamarin prover. For example, we needed to examine various SAML implementations to identify how exactly are under-specified mechanisms implemented in practice. Also, we carefully model trust establishment steps to account for arbitrary complex system and trust relationships. Finally, we develop a novel method of modelling TLS communication with Tamarin prover since we found existing approaches unsuitable for this purpose.
@@
  • We give automated formal proofs of the most important security properties of the eight protocol variants including multiple authentication, secrecy and freshness properties. The properties hold in a general setting with an unbounded number of protocol participants and sessions, and a Dolev-Yao adversary that can corrupt arbitrary participants.
  • We identify a weakness in all protocol variants — the artifact-session confusion attack. While the discovered attack does not seem to have an impact on the security of deployed systems, it highlights the necessity of closely following the SAML assertion processing rules.
  • We include the Tamarin code and the scripts needed to reproduce the results as a submission artifact and plan to make the complete model publicly available upon paper publication.
@@
Brief description of the SAML standard and the SAML Web SSO Profile is given in Section II. Then, we present the methodology and the formal model of the POST/Artifact use case in Section III. In Section IV, we present the security properties and the result of their verification with the discussion summarising the results in Section V. We discuss related work is Section VI and conclude in Section VII.
II Background
Figure 1: An example protocol execution of SAML V2.0 SSO SP-Initiated SSO with POST/Artifact bindings use case.
II-A SAML Framework
The Security Assertion Markup Language version 2.0 (in the rest of paper abbreviation SAML will denote SAML version 2.0) is a technical framework for exchanging security information between partners [23]. It is an open standard developed by the Security Services Technical Committee under OASIS. The brief technical overview of the SAML standard is given in the official document [23], whereas the standard itself is defined in a series of technical documents [21, 24, 25].
SAML framework defines the following basic building blocks: assertions, protocols, bindings and profiles. Assertions define the syntax, semantics and processing rules for security statements. Protocols define syntax, semantics and high-level processing rules for requesting and delivering assertions. Specification for both assertions and protocols includes various application-level cryptographic mechanisms (e.g., digital signatures for ensuring authenticity of authentication requests, public-key encryption for protecting the confidentiality of assertion data). Bindings map the abstract protocols to specific delivery mechanisms (e.g., SOAP, HTTP Redirect, HTTP POST, artifact). The bindings specification also mandates a transport layer security mechanism — the TLS protocol — in certain situations. Finally, profiles combine SAML assertions, protocols and bindings with an additional set of rules in a system that aims to provide a specific functionality (e.g., a SSO mechanism on the web).
@@ II-A SAML Framework
II-B Web Browser SSO Profile
@@
With HTTP artifact binding the requests and/or responses can be communicated by reference using so called artifacts. In that case, another SAML profile — the Artifact Resolution Profile — is additionally used to resolve the artifact into an actual request and/or response.
II-C SP-initiated SSO with POST/Artifact Bindings
@@
An example execution flow is given in Fig. 1. End user through the web browser client attempts to access a resource at SP. SP responds with script that initiates a HTTP POST request to the IdP containing the <AuthnRequest> structure — a SAML message consisting of, amongst other data, a unique id, SP identifier and IdP identifier. This <AuthnRequest> can be accompanied with a RelayState message which holds information about end users initial request (this is more discussed in subsection III-A). IdP receives the <AuthnRequest> and tries to authenticate the end user. After successful authentication, IdP builds the SAML <Response> message, but does not deliver it via the web browser client. Instead, it creates an <Artifact> SAML message containing an unique ID associated with the specific <Response> message. This <Artifact> is delivered to the SP via the web browser using either HTTP Redirect or POST mechanism. After receiving the <Artifact>, SP builds the <ArtifactResolve> message and sends it directly using SOAP binding to IdP. After successfully verifying the <ArtifactResolve> request, IdP responds with a <Response> containing an assertion. Finally, the SP verifies the assertion and provides the requested resource to the end user.
II-D Tamarin prover
The Tamarin prover [26, 27, 28] is state-of-the-art tool for symbolic modeling and automated analysis of security protocols. It was successfully used to analyse, break, and improve many real-world security protocols such as TLS 1.3 [29], authentication protocols for 5G networks [30], and WPA2 [31].
@@ II-D Tamarin prover
Security properties are most often specified using guarded first-order logic formulas involving actions facts, that are interpreted over execution traces. Tamarin also supports equivalence properties, but we use trace properties exclusively in this paper. The verification procedure tries to automatically verify or falsify the formula in a constraint system containing user-defined rules, built-in adversary rules and model restrictions. The automated procedure can be further refined by the user by supplying a proof-guiding oracle — a sort of heuristics for optimizing automated proving. The property, if verified, holds in a general setting with an unbounded number of protocol participants and sessions unless, of course, different assumptions are explicitly build into the model or the properties.
III Formal model
Now, we present the formal model of the POST/Artifact use case as defined in  [25, 23] (code available at [32]). To obtain the formal model of the system, we need to model the behaviour of the participants defined by the standards — namely the SPs and IdPs, but also the end users participating in the protocol by using a web browser client. Furthermore, we need to explicitly model the trust establishment steps between protocol participants that are implicit in the SSO use case. Since the use case assumes TLS connections which are not provided by default in the Tamarin prover, we build primitives that model web request-response communication over one-way authenticated TLS. Finally, we need to model the adversary that, besides the usual Dolev-Yao capabilities, can corrupt other protocol participants (SPs, IdPs, and clients) and communicate over TLS with honest participants.
III-A Methodology and Model Variants
As mentioned before, there are several degrees of freedom in the SAML specifications, and here we describe the methodology by which we converted the specifications to the the core formal model of SAML and sell as the eight protocol variants we consider in our analysis.
Our approach for optional features was as follows:
  • All required fields and processing rules (i.e., those described with MUST key word in the standard) are included in the core model, with he following exception: If the lack of processing rules gives strong evidence that the field plays a purely functional rather than security role, we leave it out of the model.
Example required fields not included in the model are the <Version> and the <Timestamp> of the <Assertion>. An example sensitive optional feature is the use of <EncryptedAssertion> instead of <Assertion> in authentication responses. This feature is clearly security relevant and we recognize as such. Note, however, we leave the implementation for future work rather than adding it to the core model right now. Table I contains the exhaustive list of all fields of <AuthnRequest> and <Response> not included in our core model with the fields we consider security sensitive.

### v2

@@ Abstract
Single Sign-On (SSO) protocols streamline user authentication with a unified login for multiple online services, improving usability and security. One of the most common SSO protocol frameworks — the Security Assertion Markup Language V2.0 (SAML) Web SSO Profile — has been in use for more than two decades, primarily in government, education and enterprise environments. Despite its mission-critical nature, only certain deployments and configurations of the Web SSO Profile have been formally analyzed. This paper attempts to bridge this gap by performing a comprehensive formal security analysis of the SAML V2.0 SP-initiated SSO with POST/Artifact Bindings use case. Rather than focusing on a specific deployment and configuration, we closely follow the specification with the goal of capturing many different deployments allowed by the standard. Modeling and analysis are performed using Tamarin prover — state-of-the-art tool for automated verification of security protocols in the symbolic model of cryptography. Technically, we build a meta-model of the use case that we instantiate to eight different protocol variants. Using the Tamarin prover, we formally verify a number of critical security properties for those protocol variants, while identifying certain drawbacks and potential vulnerabilities.
@@
security protocols, formal proof, formal verification, SAML, security analysis, Single Sign-On, SSO, Tamarin prover
@@ I Introduction
Increasing number of online services are requiring user authentication, with passwords still being the prevalent authentication method. Accessibility issues related to unique strong passwords naturally lead to password reuse, which, in turn, makes services vulnerable to various attacks [1, 2], degrading the overall security on the Internet. Simultaneously, the burden on service providers to deploy, maintain and safeguard the authentication infrastructure further compounds the complexity. In response to these challenges, the concept of Web Single Sign-On (SSO) emerged at the start of this millennium. SSO protocols enable exchange of authentication and authorization data between trusted identity providers (IdPs) and service providers (SPs), enabling users to authenticate once at an IdP and access services across multiple SPs.
@@
SSO frameworks ease the burden of authentication on SPs, but also impose stricter security requirements for IdPs and the SSO framework itself, as they are a single point of failure — a compromised IdP or a weakness in the SSO system can lead to compromises of sensitive user data across different SPs. For example, in the Sunburst supply chain attacks [5, 6] , after the initial compromise, the attackers obtained the trusted signing certificate used for signing SAML assertions, and used it to access all services within the targeted enterprise.
@@
There have been numerous efforts in empirical security analysis of OAuth 2.0 [7, 8, 9, 10, 11, 12], OIDC [13] and SAML [14, 15, 16, 17], resulting in improved protocol implementations and specifications. Whilst the empirical security analysis can identify flaws and improve security, it can not guarantee the absence of weaknesses. On the other hand, formal security analysis can offer valid proofs, sometimes machine verified or even machine generated, that the mathematical model of the protocol holds desired security properties. Of course, the security guarantees provided by a deployed system depend on the faithfulness of the mathematical model to the specification and implementation.
Formal analysis of security protocols usually employs either the symbolic or computational approach. Symbolic analysis treats cryptographic primitives as abstract operations and focuses on the logical flow of information within the protocol. On the other hand, computational model more closely resembles the real world — the messages are bitstrings, the adversary is capable of arbitrary computation (while using reasonable resources), and cryptographic primitives come with precise complexity-theoretic security guarantees. Symbolic analysis is more automation-friendly and hence more suitable for complex protocols. However, protocol proved secure in the symbolic model might still be vulnerable to real world attacks. While there are tools for automated analysis in the computational model [18, 19], the analysis is significantly more complex and still infeasible for larger protocols.
This research uses the Tamarin prover [20], a state-of-the-art tool for automated symbolic protocol analysis and — alongside ProVerif [21] — one of the most prominent tools in the field. Tamarin is well suited for modelling and analysing intricate security protocols due to its support for complex protocol flows, branching, loops, and mutable global state. These features are essential for modeling SAML V2.0 participants and their interactions, as well as the trust establishment steps. Finally, the choice of Tamarin was also influenced by the authors’ prior experience with the tool; however, we see no fundamental obstacles preventing other tools, such as ProVerif [21], from being used to comprehensively model and analyze SAML V2.0 as well.
Our research aims to conduct a systematic symbolic formal analysis and verification of the security properties associated with the SAML V2.0 SSO Profile. Hence, we would ideally formally verify that SAML SSO implementations adhering to the technical specifications outlined within the official SAML V2.0 standard documentation [22, 23] do achieve the intended security properties: Secrecy, Authentication, Authenticity and Freshness. Of course, whenever the security properties are not achieved, our goal is to identify and evaluate attacks that break the said properties, and to suggest mitigations and protocol improvements.
@@
Despite its mission-critical nature and wide adoption, SAML V2.0 has received very little attention from the formal methods’ community, with the work by Armando et al. [24, 25, 26] being the only attempt to formally analyze systems based on the SAML framework. Their research, however, is exclusively focused on the so-called Redirect/POST use case which is only one of several patterns for building Web SSO with SAML (albeit the most common one in use today). Furthermore, Armando et al. only considered one specific deployment configuration used by Google, while SAML standard allows for multiple protocol variants and features. There are no attempts at formal analysis of other use cases, let alone to model the entire specification.
@@
One of the obstacles for formal analysis is that SAML is not itself a security protocol — it is a set of building blocks that can be combined, configured and implemented in many different ways to build the protocol that achieves the desired SSO functionality. Even when we fix the specific SSO use case (e.g., the Redirect/POST use case in this paper), there are at least two additional sources of flexibility — optional features and under-specified mechanisms. While such flexibility is sometimes purely functional, there are cases where both optional features and under specified mechanisms do have an influence on the resulting security properties.
While the SP-initiated Redirect/POST binding is more common, in this paper we focus on the POST/Artifact use case. This is the first analysis of an important and widely used SAML Web SSO SP-initiated POST/Artifact use case (three EU countries mandate support for this use case as part of their eID services [27]). Moreover, SAML V2.0 documentation positions POST/Artifact use case as an alternative to Redirect/POST use case [28]. Despite its lower adoption rate due to implementation complexity on both the SP and IdP sides, the assumption is that POST/Artifact should offer potential security benefits since the authentication response is not directly transmitted within the user’s browser redirect.
@@
In this paper, we are not trying to model a specific implementation — we attempt to model the specification, accounting for the flexibility allowed by the standard. We develop a straightforward methodology that, first, identifies sensitive optional features and sensitive under-specified mechanisms. Such features and mechanisms can be instantiated in several ways, with each instantiation resulting in what we call a protocol variant. An example of optional features are digital signatures of authentication requests, which are recommended (but not required) by the SAML standard. Furthermore, we perform automated analysis of protocol variants, yielding security proofs or falsifications of desired properties. Finally, the result of the formal analysis validate the methodology and demonstrate that protocol variants do indeed have different security features.
@@
There are multiple benefits to this approach. Firstly, we can use automated formal analysis tools to examine the exact impact of each feature on the overall security guarantees provided by the protocol, as SAML standards only gives vague guidelines. Using the example of optional signatures, the standard [29] says “The SAML request MAY be signed, which provides both authentication of the requester and message integrity.” — it is unclear what exactly (if anything) can go wrong in systems configured not to require signed authentication requests. Our analysis gives an answer to this question — for the POST/Artifact use case, adding signatures to authentication requests does not have a significant influence on the security of the use case in the symbolic model.
@@
The analysis presented in this paper, while significantly more comprehensive than the related work [24, 25, 26], still falls short of fully modelling the use case — we only consider three degrees of freedom, and model only two options for each. Even with automated formal analysis, we are still unable to analyze all possible protocol variants due to the complexity explosion.
@@
In order to perform the analysis, we needed to overcome a number of technical challenges related both to SAML itself and modeling with the Tamarin prover. For example, we needed to examine various SAML implementations to identify how exactly are under-specified mechanisms implemented in practice. Also, we carefully model trust establishment steps to account for arbitrary complex system and trust relationships. Finally, we develop a novel method of modelling TLS communication with Tamarin prover since we found existing approaches unsuitable for this purpose.
Contributions
@@
  • We give automated formal proofs of the most important security properties of the eight protocol variants, including multiple authentication, secrecy and freshness properties. The properties hold in a general setting with an unbounded number of protocol participants and sessions, and a Dolev-Yao [30] adversary that can compromise arbitrary participants.
  • We identify a weakness in all protocol variants — the artifact-session confusion attack. While the discovered attack does not seem to have an impact on the security of deployed systems, it highlights the necessity of closely following the SAML assertion processing rules. We suggest a simple fix, and formally verify that it does indeed mitigate the discovered attack.
  • We include the Tamarin code and the scripts [31] needed to reproduce the results as a submission artifact, and plan to make the complete model publicly available upon paper publication.
@@
First, we discuss related work in Section II. A brief description of the SAML standard and the SAML Web SSO Profile is given in Section III. Then, we present the methodology and the formal model of the POST/Artifact use case in Section IV. In Section V, we present the security properties and the result of their verification, with the discussion summarizing the results in Section VI. Finally, we conclude in Section VII.
II Related work
The most notable example of a formal analysis of a SAML SSO use case is the research of Armando et al. [24, 25, 26]. They model and analyze the Redirect/POST use case, as well as an actual deployment by Google, using a SAT-based model checker SATMC [32]. They consider two security properties — authentication based on non-injective agreement and the resource secrecy. The most significant contribution is the discovery of a crucial flaw in the Google’s deployment — the SP identifier was left out from the <AuthnRequest> message, leading to an attack breaking both security properties considered. The authors also note that the tool fails to find similar attacks when the SP identifier is included in the message, as required by the SAML specifications.
In the only other relevant attempt at a formal analysis of SAML use cases [33], Ferdous and Poet analyze the same use case as in [26], and create a simpler model, but one that includes an explicit user authentication step and a rudimentary digital identity model. The security properties are again similar to [26] and include basic authentication and secrecy properties. They use the High-Level Protocol Specification Language (HLPSL) to specify the protocols and the Automated Validation of Internet Security Protocols and Applications (AVISPA) tool with the On-the-fly Model-Checker (OFMC) backend for analysis. Besides the SAML use case, they also investigate other SSO protocols, namely OpenID Connect and OAuth. The analysis by the tool discovered no attacks, but the authors do not make explicit claims on the exact setting in which the properties hold.
Figure 1: An example protocol execution of SAML V2.0 SSO SP-Initiated SSO with POST/Artifact bindings use case. Steps outside the scope of SAML Web Browser SSO Profile are grayed out.
III Background
III-A SAML Framework
The Security Assertion Markup Language version 2.0 (in the rest of paper, the abbreviation SAML will denote SAML version 2.0) is a technical framework for exchanging security information between partners [28]. It is an open standard developed by the Security Services Technical Committee under OASIS. The brief technical overview of the SAML standard is given in the official document [28], whereas the standard itself is defined in a series of technical documents [29, 38, 22].
SAML framework defines the following basic building blocks: assertions, protocols, bindings and profiles. Assertions define the syntax, semantics and processing rules for security statements. Protocols define syntax, semantics and high-level processing rules for requesting and delivering assertions. Specification for both assertions and protocols includes various application-level cryptographic mechanisms (e.g., digital signatures for ensuring authenticity of authentication requests, public-key encryption for protecting the confidentiality of assertion data). Bindings map the abstract protocols to specific delivery mechanisms (e.g., SOAP, HTTP Redirect, HTTP POST, artifact). The bindings’ specification also mandates a transport layer security mechanism — the TLS protocol — in certain situations. Finally, profiles combine SAML assertions, protocols and bindings with an additional set of rules in a system that aims to provide a specific functionality (e.g., a SSO mechanism on the web).
@@ III-A SAML Framework
Message Relevant contents Protocol step description
Resource request Resource URL User attempts to access resource on SP, no valid session exists. SP stores the requested resource URL using the RelayState mechanism.
Authentication Request IdP, RelayState, AuthnRequest(OptionalSignature, id_authnrequest, SP, IdP) SP sends user an HTML form containing encoded SAML AuthnRequest (SAMLRequest) and RelayState. User submits form or script auto-submits to IdP’s SSO service.
Request for user authentication Authentication challenge IdP checks for existing session or prompts user for credentials.
User authenticates User credentials IdP verifies user credentials (if needed) and establishes session.
Artifact RelayState, Artifact(messageHandle) IdP creates a SAML Response containing user assertion. IdP creates an artifact referencing the Response and redirects user’s browser with the artifact.
Artifact Resolve ArtifactResolve(id_artifactresolve, SP, messageHandle, ArtifactResolveSignature) SP sends the received Artifact to the IdP’s Artifact Resolution Service to retrieve the original SAML Response.
Response ArtifactResponse( IdP, id_artifactresolve,    Response( ResponseSignature, id_response, IdP, id_arthrequest,     Assertion( id_assertion, IdP, Audience(SP), assertion_secret,      Subject(clientID, id_authnrequest) ) ) ) The IdP sends to the SP the SAML Response referenced by received artifact.
Resource Resource The SP extracts the SAML Response from the ArtifactResponse validates it (if not already signed), and processes the embedded assertion. Based on the assertion, the SP creates a local security context for the user. Finally, the SP retrieves the originally requested resource URL from the RelayState and redirects the user’s browser to access that resource.
Table I: Description of SAML POST/Artifact messages from a typical execution flow.
III-B Web Browser SSO Profile
@@
With HTTP artifact binding the requests and/or responses can be communicated by reference using so-called artifacts. In that case, another SAML profile — the Artifact Resolution Profile — is additionally used to resolve the artifact into an actual request and/or response.
III-C SP-initiated SSO with POST/Artifact Bindings
@@
A high level execution flow is given in Fig. 1, whilst each message contents are given in Table I. End user through the web browser client attempts to access a resource at SP. SP responds with a script that initiates an HTTP POST request to the IdP containing the <AuthnRequest> structure — a SAML message consisting of, amongst other data, a unique id, SP identifier and IdP identifier. This <AuthnRequest> can be accompanied by a RelayState message which holds information about end user’s initial request (this is more discussed in Section IV-A). IdP receives the <AuthnRequest> and tries to authenticate the end user. After successful authentication, IdP builds the SAML <Response> message, but does not deliver it via the web browser client. Instead, it creates an <Artifact> SAML message containing an unique ID associated with the specific <Response> message. This <Artifact> is delivered to the SP via the web browser using either HTTP Redirect or POST mechanism. After receiving the <Artifact>, SP builds the <ArtifactResolve> message and sends it directly using SOAP binding to IdP. After successfully verifying the <ArtifactResolve> request, IdP responds with a <Response> containing an assertion. Finally, the SP verifies the assertion and provides the requested resource to the end user.
III-D Tamarin prover
The Tamarin prover [20, 39, 40] is a state-of-the-art tool for symbolic modeling and automated analysis of security protocols. It was successfully used to analyze, break, and improve many real-world security protocols such as TLS 1.3 [41], authentication protocols for 5G networks [42], and WPA2 [43].
@@ III-D Tamarin prover
Security properties are most often specified using guarded first-order logic formulas involving actions facts, that are interpreted over execution traces. Tamarin also supports equivalence properties, but we use trace properties exclusively in this paper. The verification procedure tries to automatically verify or falsify the formula in a constraint system containing user-defined rules, built-in adversary rules and model restrictions. The automated procedure can be further refined by the user by supplying a proof-guiding oracle — a sort of heuristics for optimizing automated proving. The property, if verified, holds in a general setting with an unbounded number of protocol participants and sessions unless, of course, different assumptions are explicitly built into the model or the properties.
IV Formal model
Now, we present the formal model of the POST/Artifact use case as defined in [22, 28]. To obtain the formal model of the system, we need to model the behaviour of the participants defined by the standards — namely the SPs and IdPs, but also the end users participating in the protocol by using a web browser client. Furthermore, we need to explicitly model the trust establishment steps between protocol participants that are implicit in the SSO use case. Since the use case assumes TLS connections, which are not provided by default in the Tamarin prover, we build primitives that model web request-response communication over one-way authenticated TLS. Finally, we need to model the adversary that, besides the usual Dolev-Yao capabilities, can compromise other protocol participants (SPs, IdPs, and clients) and communicate over TLS with honest participants.
IV-A Methodology and Model Variants